John deposits $450 at the end of every 3 months in a savings account. The account has a 5% interest compounded quarterly. How much will he have in his account at the end of 7 years and 3 months?


1
Expert's answer
2022-02-25T13:10:42-0500

The amount at the end of 7 years and 3 months is given by:

"FV= \frac{PV ((1+\frac{r}{m})^{nm}-1)}{(\frac{r}{m})}"


Where FV is the future amount in the account = ?

PV is the amount deposited every three months = $450

r is the rate of interest on the account = 5%

n is the number of years = 7 years and 3 months (7.25)

m is the frequency of compounding = 4


"\therefore FV=\frac{450(1+\frac{0.05}{4})^{(7.25×4)}-1)}{(\frac{0.05}{4})}"

"FV=\$15612.92"